OpenStudy (anonymous):

given the following exponential function below, what is the value of f(1/2)? (round your answer to the nearest hundredth.) f(x)=-e^2x-3

OpenStudy (anonymous):

anyone? pleasee :((((

OpenStudy (anonymous):

.14

OpenStudy (anonymous):

thats not one of the answers i have

OpenStudy (anonymous):

Is this the equation?\[f(x)=e ^{2x}-3\] If so, then f(1/2) looks like this\[f(1/2)=e ^{2*\frac{ 1 }{ 2}}-3\]\[f(1/2)=e ^{1}-3\]\[f(1/2)=2.72-3 = -.28\]